OpenStudy (anonymous):

twice the difference of a number and 3 is the same as 1 added to three times the number

OpenStudy (zzr0ck3r):

twice the the difference of a number and 3 is 2|x-3| 1 added to three times the number 1+3x so 2|x-3|=1+3x solve for x

OpenStudy (anonymous):

Thank you! I thought it was 2(x-3)=3(x+1) but I kept getting the wrong answer

OpenStudy (zzr0ck3r):

you shuold have two answers

OpenStudy (zzr0ck3r):

should*

OpenStudy (zzr0ck3r):

this is actually quite ambiguous, if they only want one answer use 2(x-3) = 1+ 3x

OpenStudy (anonymous):

the convention is that whichever is said first is assumed to be bigger. At this math level, it's to much trouble to introduce absolute value so it's really 2(n-3) 1 = 1 + 3n

OpenStudy (zzr0ck3r):

but the difference between a and b is |a-b|=|b-a|, but sometimes when someone says the difference of a and b, they mean a-b (the first thing they say minus the second), but technically the difference of two numbers is always positive and we do that by |a-b|

OpenStudy (zzr0ck3r):

convention ......I use definition :)

OpenStudy (anonymous):

Exactly, thank you! I solved for x and got x=-7

OpenStudy (zzr0ck3r):

the difference of a, and b should be the same as the difference of b, and a. so this convention is silly :(

OpenStudy (anonymous):

For those kids who are just learning basic algebra, it's really difficult to explain what absolute value really means. They're only taught that absolute value simply means positive. So to avoid any confusion, whenever they make up these questions that has difference involved, they will ways make sure the bigger number always comes first

OpenStudy (anonymous):

they will *always* make sure

OpenStudy (zzr0ck3r):

this is not good enough justification to me:) tell anyone to look at a number line and give the distance between -5 and 3 and they will say 8. But it is what it is.

OpenStudy (anonymous):

@sourwing can you help me?? when ur done here

OpenStudy (anonymous):

actually that's really how it is. Kids are taught to subtract small number from bigger number first. So when ask them for the difference, they will always do big number - small numbers